Three identical particles each of mass m are placed at the corners of an equilateral triangle of side l.The work done by external force to increase the side of triangle from l to 2l is


a) -3GM^2/2l


b) -3Gm^2/l


c) 3GM^2/2l


d) 3GM^2/l

Grade:11

2 Answers

Swapnil Saxena
12 years ago

The gravitational potential energy of the a pair of particles is -Gmm/l2

Since there exist three pairs (3 particles constitute three pairs)  then the potential energie of system is -3(Gmm/l)

When the side of the cube is increased frm l to 2l then the potential energy of system is   -3(Gmm/(2l))

Subtracting final Uf with initial Ui = 3(Gmm/(2l))

Note: If we are bringing the particles away frm each other then we have to supply energy against gravitation so it is definitely +. So no chance for option a,b

So clearly Option (c)

Ap
5 years ago
If U= uf-ui
    U= -GMm/r
Wd=3Gm^2/2l
Hense option 3 is correct.
In this way you can solve other given problems .